How to solve greatest common divisor ( g.c.d.) of 1132 and 164

Grade:12th pass

4 Answers

Aditya Gupta
4 years ago
We first perform the prime factorisation:
1132= 2^2*283
And 164= 2^2*41
Clearly, the common factors are 2^2.
So the gcd is equal to 2^2= 4.
KINDLY APPROVE :))
Saurabh Koranglekar
4 years ago
Dear student

To understand the process

Please find the link below

https://www.askiitians.com/revision-notes/class-6-maths/playing-with-numbers/

Hcf is 4

Regards
Vikas TU
4 years ago
We found the factors and prime factorization of 160 and 164. The biggest common factor number is the GCF number. So the greatest common factor 160 and 164 is 4.
Yashraj
3 years ago
this is a easy question i think you should revis e once 
just make factors and find common part
thanks